# A Model-Consistent Data-Driven Computational Strategy for PDE Joint Inversion Problems

### v2

@@ 5 Impact of learning inaccuracy on reconstruction
While numerical examples in rather academic environments show promises in learning relations between the coefficients from given data, it is not possible to learn the exact relation between the unknowns in practice, even if such an exact relation exists. In other words, the map $\mathcal{N}_{\widehat{\theta}}$ that we learned is only an approximation to the true relation. In this section, we show that, under reasonable assumptions on the inverse problems, when the learning error is under control, the error it caused in the joint reconstruction is also under control.
@@
For the simplicity of presentation, we assume that the domain $\Omega$ is bounded with smooth boundary $\partial\Omega$.
@@ Acknowledgments
We would like to thank Yan Cheng for his help on the numerical experiment in Section 4.2. This work is partially supported by the National Science Foundation through grants DMS-1913309, DMS-1937254, and EAR-2000850.

### v3

@@ 5 Impact of learning inaccuracy on reconstruction
While numerical examples in rather academic environments show promises in learning relations between the coefficients from given data, it is not possible to learn the exact relation between the unknowns in practice, even if such an exact relation exists. In other words, the map $\mathcal{N}_{\widehat{\theta}}$ that we learned is only an approximation to the true relation. In this section, we show that, under reasonable assumptions on the inverse problems, when the learning error is under control, the error caused in the joint reconstruction is also under control.
@@
For the simplicity of presentation, we assume that the domain $\Omega$ is bounded with a smooth boundary $\partial\Omega$.
@@ 5.1 Joint inversion of the diffusion model
One should also notice that the above characterization of error propagation is relative in the sense that the quantity $\frac{H}{\sigma}$ appears on the right-hand side of the inequality. In other words, the stability estimate does not directly imply that we can reconstruct $\sigma$ accurately. In fact, it only says that if we have a reconstruction algorithm, and we use the relations $\mathcal{N}_{\theta}$ and $\mathcal{N}_{\widetilde{\theta}}$ in the reconstruction process, then $\mathcal{N}_{\theta}$ and $\mathcal{N}_{\widetilde{\theta}}$ being close implies that $\sigma$ and $\widetilde{\sigma}$ are close, even though both reconstructions can be fairly inaccurate.
@@ Acknowledgments
We would like to thank the anonymous referees for their useful comments that helped us improve the quality of this work. We would like to thank Yan Cheng for his help on the numerical experiment in Section 4.2. This work is partially supported by the National Science Foundation through grants DMS-1913309, DMS-1937254, and EAR-2000850.